OpenStudy (bloomlocke367):

A student deposits the same amount of money into her bank account each week. At the end of the second week she has $30 in her account. At the end of the third week she has $45 in her account. How much will she have in her bank account at the end of the ninth week?

OpenStudy (bloomlocke367):

Should I just write an explicit formula?

OpenStudy (igreen):

Week 2 = 30 Week 3 = 45 I guess we can make an equation.. (2, 30) and (3, 45) Plug them into the slope formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{45-30}{3-2}\) Simplify

OpenStudy (bloomlocke367):

I'm assuming that at the end of the first week she had $15... should I make that assumption? and why did you make that equation? Would it be wrong to use:\[a _{n}=a _{1}+(n-1)d\]

OpenStudy (igreen):

No..it wouldn't.

OpenStudy (bloomlocke367):

Okay. Which is easier?

OpenStudy (igreen):

Well, just use the formula since that's what they want you to do.

OpenStudy (igreen):

'a' is the first term, which you are correct, is 15. And 'd' is the common difference.

OpenStudy (igreen):

To find 'd', just subtract 45 - 30

OpenStudy (bloomlocke367):

\[a _{n}=15+(n-1)15\] \[ =15+15n-15\] \[=15n\] right?

OpenStudy (igreen):

Yep, now plug in '9' for 'n'

OpenStudy (igreen):

That will give us the amount she has on the 9th week.

OpenStudy (bloomlocke367):

135

OpenStudy (igreen):

Correct

OpenStudy (bloomlocke367):

That was easier than I thought... I saw the word problem and kinda freaked out XD

OpenStudy (igreen):

Finding the slope would've gave us 15.. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{45-30}{3-2}\) \(m = \dfrac{15}{1}\) \(m = 15\) So the equation would've been y = 15x, and we plug in '9' for 'x' and we get the same thing :P

OpenStudy (bloomlocke367):

Okay. :) thanks.